### v1

# Towards Understanding Variants of Invariant Risk Minimization through the Lens of Calibration

Abstract
Machine learning models traditionally assume that training and test data are independently and identically distributed. However, in real-world applications, the test distribution often differs from training. This problem, known as out-of-distribution generalization, challenges conventional models. Invariant Risk Minimization (IRM) emerges as a solution, aiming to identify features invariant across different environments to enhance out-of-distribution robustness. However, IRM’s complexity, particularly its bi-level optimization, has led to the development of various approximate methods. Our study investigates these approximate IRM techniques, employing the Expected Calibration Error (ECE) as a key metric. ECE, which measures the reliability of model prediction, serves as an indicator of whether models effectively capture environment-invariant features. Through a comparative analysis of datasets with distributional shifts, we observe that Information Bottleneck-based IRM, which condenses representational information, achieves a balance in improving ECE while preserving accuracy relatively. This finding is pivotal, as it demonstrates a feasible path to maintaining robustness without compromising accuracy. Nonetheless, our experiments also caution against over-regularization, which can diminish accuracy. This underscores the necessity for a systematic approach in evaluating out-of-distribution generalization metrics, one that beyond mere accuracy to address the nuanced interplay between accuracy and calibration.
1 Introduction
In the evolving landscape of machine learning, the importance of out-of-distribution (OOD) generalization (Arjovsky, 2020) and calibration (Guo et al., 2017) cannot be understated, especially in real-world applications and critical scenarios. Traditional machine learning approaches, grounded in the assumption that data are independently and identically distributed (IID), often struggle to cope with OOD scenarios. Additionally, a notable trend in contemporary machine learning models is their tendency towards overconfidence, which undermines the reliability of their confidence estimations. This issue is primarily recognized as a calibration challenge (Guo et al., 2017).
Responding to these limitations, there has been a surge in research focusing on methodologies like Invariant Risk Minimization (IRM) (Arjovsky et al., 2020). IRM is designed to identify features that remain constant across different environments, thereby facilitating more robust generalization in the face of environmental variations. However, the computational demands of IRM have led to the exploration of more feasible, approximate methods. Despite the development of several IRM variants, their inability to consistently surpass the performance of finely-tuned Empirical Risk Minimization (ERM) models has been observed (Gulrajani & Lopez-Paz, 2020)(Zhang et al., 2023). This performance gap raises critical questions about the inherent deficiencies of these IRM variants. As we showed in Figure 1, even by leveraging IRM variants, there is a trade-off between in-distribution accuracy and OOD accuracy. It implies that models trained by IRM variants failed to obtain invariant features since they sacrifice in-distribution accuracy to achieve better OOD accuracy.
From another perspective of IRM, it can also be conceptualized as a specialized form of multi-domain calibration (Wald et al., 2021). Our study intersects with these findings, examining the extent to which IRM variants mimic the original model in terms of calibration. We posit that effective implementation of IRM should lead to reduced loss and a zero expected calibration error (ECE) (Wald et al., 2021). Through comparative experimentation, we assess various IRM variants in environments marked by distributional shifts, such as diversity and correlation shifts (Ye et al., 2021). Our evaluation focuses not only on accuracy but also on ECE. The results highlight notable disparities in ECE among different methodologies, even when OOD accuracy levels are similar. This suggests that a sole reliance on OOD accuracy for method development might be inadequate. Our observations include a range of calibration behaviors across methods, from overconfidence to underestimation.
Our key contributions are as follows:
  • We investigate the reasons behind the gap between IRM formulation and its variants through the lens of ECE.
  • We demonstrate that Information Bottleneck-based IRM (IB-IRM) effectively lowers ECE, aligning it more closely with the original objectives of IRM (Figure 3, Table 2).
  • We reveal that IRM variants with robust regularization typically achieve lower ECE (Figure 3).
  • We identify a trade-off between ECE and accuracy, thereby underscoring the intricate balance and compromises necessary for achieving both robust generalization and accurate calibration (Figure 5).
Utilizing ECE as a metric, our study sheds light on the gap between the original formulation of IRM and its variants. We recognize a notable trade-off between ECE and accuracy, highlighting the complexities and necessary compromises in striving for both robust generalization and precise calibration.
ID IRMv1 OOD IRMv1
ERM IRMv1 BIRM IB-IRM IRM Game PAIR
Figure 1: A graph observed in CMNIST showing the trade-off between in-distribution (ID) and OOD accuracy in approximation methods of IRM. Figure (a) visualizes the accuracy of the IRMv1 model ID and OOD, plotted against the training steps on the horizontal axis. The two solid lines are approximately symmetric around an accuracy of 50%. Figure (b) visualizes the ID accuracy (horizontal axis) against the OOD accuracy (vertical axis) for typical approximation methods of IRM. There is a clear trend that as the accuracy improves OOD, the ID accuracy decreases.
2 Background
2.1 OOD Generalization
In traditional deep neural network scenarios, it is assumed that the learning and test environments are IID. The most commonly used learning algorithm in this assumption has been ERM(Vapnik, 1991), which has been successful in various scenarios. Specifically, ERM aims to minimize the sum of losses of the model $f:\mathcal{X}\to\mathcal{Y}$ in the training environment $E_{train}$.
$\min_{f:\mathcal{X}\to\mathcal{Y}}\sum_{e\in E_{train}}R^{e}(f)$ (1)
However, in the real world, the assumption of IID does not always hold, and distribution shifts between training and testing environments can occur (Quionero-Candela et al., 2009). When the model relies on features that are only effective in the training environment, known as shortcut features, its performance can decrease in the testing environment (Geirhos et al., 2020). This issue is commonly known as OOD generalization and has become an urgent problem in recent years. Therefore, it is necessary to train robust models that are not dependent on specific environments and can perform consistently across all environments. A robust model is defined as follows.
$\mathbb{E}[Y^{e}|f(X^{e})]=\mathbb{E}[Y^{e^{\prime}}|f(X^{e^{\prime}})],\forall e,e^{\prime}\in E_{all}$ (2)
where $E_{all}$ represents the set of all possible environments. (2) indicates that, for a robust model, the conditional probability of the label given the model’s prediction should be equal between any two environments.
Since it is difficult to access all real-world environments during training, various methods have been proposed to realize invariant models satisfying (2) using limited learning environments.
2.2 Calibration
The alignment of a model’s predictive confidence with its actual accuracy is extremely important in practical scenarios such as medical image diagnosis and autonomous driving. For example, if there are 100 instances where a model predicts with 80% confidence, then 80 of those predictions should be correct. However, in recent deep neural networks, issues such as increased model capacity have led to problems with the miscalibration and many methods have been proposed to address this calibration issue (Guo et al., 2017). The correct calibration in all environments is defined for any predictive probability $\alpha$ of $f$ as follows:
$\mathbb{E}[Y^{e}|f(X^{e})=\alpha]=\alpha,\forall e\in E_{all}$ (3)
Equation (3) implies that the model’s predictive probability is always consistent with its actual accuracy.
Furthermore, the quality of uncertainty calibration is often quantified as ECE and is commonly used. It is defined as follows:
$ECE=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|Acc(B_{m})-Conf(B_{m})|$ (4)
$Acc(B_{m})$ and $Conf(B_{m})$ are defined as follows
$Acc(B_{m})=$ $\frac{1}{\left|B_{m}\right|}\sum_{b\in B_{m}}\mathbf{1}\left(\hat{y}_{b}=y_{b}\right)$
$Conf(B_{m})=$ $\frac{1}{\left|B_{m}\right|}\sum_{b\in B_{m}}\hat{p}_{b}$
where the predictive probabilities ranging from 0 to 1 are divided into $M$ bins, and for each bin, $Acc$ represents the accuracy of predictions within the bin, while $Conf$ represents the average predictive probability (confidence level) within the bin.
2.3 OOD Generalization and Calibration
It has been pointed out that OOD generalization and calibration are interrelated. Notably, (3) can be a sufficient condition for (2), theoretically demonstrating that calibration across all environments leads to OOD generalization.
Additionally, it has been empirically demonstrated that there is a correlation between a model’s accuracy in OOD (OOD) scenarios and its calibration performance (Ovadia et al., 2019).
In this paper, we utilize this fact and hypothesize that models with high calibration performance are likely to achieve better OOD generalization. We evaluate the approximation methods of Invariant Risk Minimization (IRM), introduced in 4, based on the model’s calibration performance to assess their OOD generalization capabilities.
3 Preliminary
3.1 Domain Invariant Representation
One approach to making models more robust in OOD scenarios is through representation learning. Representation learning is a method that teaches the model to learn effective data representations from raw data for better performance in predictions. Traditionally, features of data were designed by human, but representation learning reduces this necessity and can also compress the dimensions of input data, enabling faster computations.
In the context of OOD scenarios, there is an application for Representation Learning, specifically aimed at extracting environment-invariant features from raw data and reducing environment-specific spurious features to achieve more robust predictions. For example, in a task of classifying animals from images, if the model learns to associate parts of an animal with extraneous elements, it may lead to high performance within the training domain but fail to make robust predictions when the background distribution shifts in the test domain. However, if the model discards spurious features (e.g., background) from the raw data, it can consistently make robust predictions across all environments. One particularly well-known method that applies representation learning to out-of-distribution generalization scenarios is Domain-Adversarial Neural Networks (DANN) (Ganin et al., 2016). This approach uses adversarial training to adjust the learning process so that it is impossible to discern from the extracted features which environment the input comes from, thereby aiming to extract invariant features.
3.2 Invariant Learning
In the context of applying representation learning to OOD generalization, the method we focus on in this paper is IRM(Arjovsky et al., 2020). IRM is proposed with the aim of extracting environment-invariant features from input data to enable consistent predictions across any environment. IRM is defined with the purpose of realizing (2) as follows.
$\min_{\begin{subarray}{c}\Phi:\mathcal{X}\to\widehat{\mathcal{H}}\\ \omega:\widehat{\mathcal{H}}\to\widehat{\mathcal{Y}}\end{subarray}}\sum_{e\in E_{train}}R^{e}(\omega\circ\Phi)$ (5)
$subject\ to\ \ \ \omega\in\arg\min_{\bar{\omega}:\widehat{\mathcal{H}}\to\widehat{\mathcal{Y}}}R^{e}(\bar{\omega}\circ\Phi),\forall e\in E_{train}$
Here, $\mathcal{X}$ represents the input space of each data point, $\widehat{\mathcal{H}}$ denotes the extracted invariant feature space, and $\mathcal{Y}$ is the output space of the model. IRM defines the function as $f=\omega\circ\Phi$, where $\Phi$ maps the input data to $\widehat{\mathcal{H}}$, and based on these extracted invariant features, the final prediction is made by $\omega$ The aim of IRM is to achieve consistent predictions as a form of OOD generalization.
The formulation in (5) represents a bi-level optimization problem, which is challenging to solve exactly. As a result, various implementable approximation methods have been proposed. However, despite numerous contributions to the development of approximation methods for IRM, a method that generally surpasses well-tuned ERM in terms of out-of-distribution generalization accuracy has not yet been proposed. Therefore, it is essential to understand why these approximation methods have not succeeded, and what approaches might be more suitable for addressing the approximation problem.
4 Related Works
Table 1 presents an overview of IRM and its approximation methods. We discuss the variants of IRM in the following section.
Methods Featurizer Predictor Linearity of $\omega$ Information Bottleneck Game theory Ensemble Bayesian inference
IRM $\Phi$ $\omega$
IRMv1 $\omega\cdot\Phi$ 1.0
IB-IRM $\omega\cdot\Phi$ 1.0
PAIR $\omega\cdot\Phi$ 1.0 (Pareto Optimality)
[…]
BIRM $\omega\cdot\Phi$ 1.0
Table 1: Approximation methods of IRM
4.1 IRMv1
IRMv1(Arjovsky et al., 2020) constrains $\omega$ in (5) as a linear mapping and allows the transformation $\Phi^{\prime}=(\omega\cdot\Phi)$, $\omega^{\prime}=1.0$. The following one-variable optimization problem can be relaxed and computed.
$\min_{\Phi^{\prime}:\mathcal{X}\to\mathcal{Y}}\sum_{e\in E_{train}}R^{e}(\Phi^{\prime})\ +\ \lambda\parallel\nabla_{\omega^{\prime}\mid\omega^{\prime}=1.0}R^{e}(\omega^{\prime}\cdot\Phi^{\prime})\parallel^{2}$ (6)
4.2 Information Bottleneck based IRM(IB-IRM)
While IRMv1 has been successful with certain OOD generalization datasets such as CMNIST, it has been noted that if not enough invariant features are included in each environment in the training data, robust prediction is lost. Therefore, IB-IRM(Ahuja et al., 2022) aims to improve IRMv1 by compressing the entropy of $\Phi$ using the information bottleneck method, thereby preventing excessive reliance on features that adversely affect the invariant prediction. The variance of the features extracted from each data by $\Phi$ is added as a regularization term.
$\lambda\parallel\nabla_{\omega^{\prime}\mid\omega^{\prime}=1.0}R^{e}(\omega^{\prime}\cdot\Phi^{\prime})\parallel^{2}+\ \gamma Var(\Phi)$ (7)
4.3 Pareto IRM(PAIR)
In general, there is a trade-off between ERM and OOD generalization, and PAIR(Chen et al., 2023) focuses on the need to properly manage that trade-off when generalizing a model out of distribution. To achieve more robust models, PAIR is designed to find their Pareto optimal solutions in terms of multi-objective optimization. In fact, it is implemented as a multi-objective optimal problem for ERM, IRMv1 penalties, and vREx(Krueger et al., 2021) penalties.
$\min_{\Phi^{\prime}:\mathcal{X}\to\mathcal{Y}}(\mathcal{L}_{ERM},\mathcal{L}_{IRMv1},\mathcal{L}_{vREx})$ (8)
where
$\mathcal{L}_{ERM}$ $=\sum_{e\in E_{train}}R^{e}(\Phi^{\prime})$
$\mathcal{L}_{IRMv1}$ $=\parallel\nabla_{\omega^{\prime}\mid\omega^{\prime}=1.0}R^{e}(\omega^{\prime}\cdot\Phi^{\prime})\parallel^{2}$
$\mathcal{L}_{vREx}$ $=Var(\{R^{e}(\Phi^{\prime})\}_{e\in E_{train}})$
4.4 IRM Game
IRM Game(Ahuja et al., 2020) aims to achieve Nash equilibrium in terms of inference accuracy by introducing a game-theoretic framework to the IRM between each environment. Each environment $e$ in the training data is assigned its own predictor $\omega^{e}$, each of which is trained to be optimal in each environment, and the final $\omega_{ens}$ is the ensemble of all predictors $\frac{1}{|E_{train}|}\sum_{e\in E_{train}}\omega^{e}$. By learning environment-specific predictors, the limitation on linearity made in IRMv1 is eliminated, and an implementation closer to (5) is aimed at.
$\min_{\begin{subarray}{c}\Phi:\mathcal{X}\to\widehat{\mathcal{H}}\\ \omega_{ens}:\widehat{\mathcal{H}}\to\widehat{\mathcal{Y}}\end{subarray}}\sum_{e\in E_{train}}R^{e}(\omega_{ens}\circ\Phi)$ (9)
$s.t.\ \omega^{e}\in\arg\min_{\bar{\omega}^{e}:\widehat{\mathcal{H}}\to\widehat{\mathcal{Y}}}R^{e}\biggl{\{}\bar{\omega}^{e}_{ens}\circ\Phi\biggr{\}},\forall e\in E_{train}$
where $\bar{\omega}^{e}_{ens}=\frac{1}{|E_{train}|}\Bigl{(}\bar{\omega}^{e}+\sum_{e^{\prime}\neq e}w^{e^{\prime}}\Bigr{)}$
4.5 Bayesian IRM(BIRM)
In BIRM(Lin et al., 2022), IRMv1 may overfit the training environment if the training data are insufficient, so we introduced a Bayesian estimation approach there. Noting that the posterior probability $p(\omega^{e}|\Phi(X^{e}),Y^{e})$ is invariant in all environments given the data mapped by $\Phi$ and the correct label if the model is able to acquire invariant features, the penalty of IRMv1 is modified. By using Bayesian estimation instead of point estimation, they aim to improve the generalization performance of the model by taking into account the uncertainty of the data.
5 Experiments
Although the approximation methods in the previous chapter have achieved some success in terms of OOD accuracy on OOD datasets (Arjovsky et al., 2020)(Ahuja et al., 2022)(Chen et al., 2023)(Ahuja et al., 2020)(Lin et al., 2022), it is not known how invariant features they actually acquire and how robust they are to unknown environments. Therefore, we evaluated them by comparing their calibration performance with ECE.
5.1 Setting
We compared ECE of each model on four different OOD-datasets. Specifically, the datasets used were Colored MNIST (CMNIST)(Arjovsky et al., 2020), Rotated MNIST (RMNIST)(Ghifary et al., 2015), PACS(Li et al., 2017), and VLCS(Fang et al., 2013), sourced from the DomainBed benchmark(Gulrajani & Lopez-Paz, 2020). These datasets can be distinguished based on the type of distribution shift they represent(Ye et al., 2022). The distribution shift in CMNIST is known as the Correlation Shift, where the conditional probability of labels given spurious features changes across environments. The other three datasets—RMNIST, PACS, and VLCS—are classified under Diversity Shift, where the prior probability of spurious features changes across environments.
For optimization, Adam was used consistently across all models, and the tuning of learning rates and hyperparameters for each method was conducted in accordance with their respective papers.
5.2 Correlation Shift
5.2.1 Dataset
CMNIST is a dataset based on MNIST, containing 70,000 examples of dimension (1, 28, 28). It is adapted for a binary classification task where digits less than 5 are labeled as class 0, and those 5 or above as class 1. Spurious correlations are added by assigning colors (red or green) to the digits, with the proportion of the two colors varying in each class depending on the environment. Denoting the proportion of green in class 0 as environment $e$, then the training environments are set as $e=[0.1,0.2]$, and the test environment as $e=[0.9]$.
5.2.2 Results
ERM IRMv1 BIRM IB-IRM IRM Game PAIR 100% PREDS 50% PREDS
Figure 2: Comparison of OOD accuracy (horizontal axis) and OOD ECE (vertical axis) in CMNIST. The red solid line represents the theoretical values when all model’s predictive probabilities are 100%, and the blue solid line represents the case when all predictive probabilities are 50%. IRMv1, IB-IRM, and BIRM, which achieve relatively high inference accuracy, are distributed near the blue line. This indicates that these models mitigate reliance on non-invariant features by making predictions more ambiguous.
(a) OOD Accuracy
(b) OOD ECE
1 10 100 1000 10000
$\lambda$ in IRMv1
Figure 3: Comparison of the model’s OOD ECE in CMNIST, with various values of $\lambda$ in the formulation of IRMv1 6. It was observed that increasing $\lambda$, thereby increasing the regularization penalty in IRMv1, results in a lower ECE for the model in OOD environment.
Fig. 2 presents a scatter plot showing the relationship between OOD accuracy (horizontal axis) and ECE (vertical axis) for various methods, each trained with multiple hyperparameters. The red solid line represents the theoretical value when all model’s predictive probabilities are 100%, indicating very high confidence, while the blue solid line represents the case when all predictive probabilities are 50%, indicating very low and ambiguous confidence. The methods with relatively high accuracy (IRMv1(green), IB-IRM(light blue), BIRM(red)) align with the blue solid line, showing that they can overcome excessive fitting to non-invariant features by making predictions more ambiguous. On the other hand, ERM(blue) and IRM Game(purple) tend to be distributed near the red solid line, indicating predictions with relatively high confidence. PAIR(yellow) is distributed between the two solid lines, demonstrating characteristics that are intermediate between the two aforementioned groups.
Figure 3 visualizes how the penalty introduced in IRMv1 for the purpose of out-of-distribution generalization actually affects the ECE. It specifically compares different values of $\lambda$ in IRMv1’s formulation 6, demonstrating that as the impact of IRMv1’s penalty increases, the ECE decreases, suggesting better calibration. Additionally, the effects of the two regularization penalties in IB-IRM are similarly shown in the Appendix, Figures 6 and 7. In each case, comparisons were made by varying the magnitude of the regularization coefficient of the focused penalty while keeping the other penalty sufficiently influential. These results, like those in Figure 3, suggest that each penalty contributes to the improvement of the model’s ECE.
5.3 Diversity Shift
5.3.1 Datasets
The details for RMNIST, PACS, and VLCS are as follows:
RMNIST: RMNIST is a dataset based on MNIST, containing 70,000 examples of the same dimension (1, 28, 28) and includes 10 classes. The environments are distinguished by the rotation angle of the digits, denoted as $e$. The training environments are set at $e=[15^{\circ},30^{\circ},45^{\circ},60^{\circ},75^{\circ}]$, while the test environment is at $e=[0^{\circ}]$.
PACS: PACS is a dataset containing 9,991 examples with dimensions (3, 224, 224) and includes 7 classes. The environments are distinguished by four different styles of images. Specifically, the training environments are set as $e=[Cartoons,Photos,Sketches]$, and the test environment is set as $e=[Art]$.
VLCS: VLCS is a dataset containing 10,729 examples with dimensions (3, 224, 224) and includes 5 classes. There are four types of environments. The training environments are set as $e=[Caltech101,LabelMe,SUN09]$, and the test environment is set as $e=[VOC2007]$.
5.3.2 Results
(a) RMNIST
(b) PACS
(c) VLCS
ERM IRMv1 BIRM IB-IRM IRM Game PAIR ID-ECE = OOD-ECE
Figure 4: Comparison of the relationship between ECE in the training environment (horizontal axis) and the test environment (vertical axis). The red solid line represents the case where the ECE is equal in both environments. It was observed that IB-IRM (in light blue) is distributed near the red solid line, indicating a tendency not to overfit to the training environment compared to other methods.
ERM IRMv1 IB-IRM BIRM IRM Game PAIR
RMNIST ACC 91.1$\pm$1.0 91.9$\pm$1.0 90.9$\pm$0.9 89.1$\pm$0.4 90.8$\pm$1.0 87.4$\pm$0.8
ECE 5.27$\pm$1.11 3.98$\pm$0.69 1.58$\pm$0.44 6.93$\pm$0.56 7.17$\pm$0.98 7.88$\pm$0.79
PACS ACC 73.6$\pm$2.0 74.9$\pm$1.9 75.9$\pm$0.9 71.5$\pm$1.7 73.2$\pm$1.3 76.3$\pm$1.2
ECE 11.99$\pm$1.86 13.35$\pm$1.45 12.85$\pm$0.60 13.73$\pm$1.93 12.50$\pm$1.45 12.46$\pm$0.95
VLCS ACC 70.3$\pm$1.0 70.5$\pm$1.2 69.4$\pm$0.7 68.8$\pm$0.8 69.9$\pm$1.2 71.1$\pm$1.2
ECE 12.53$\pm$1.42 11.90$\pm$1.26 7.84$\pm$1.55 10.57$\pm$1.81 8.70$\pm$0.93 10.34$\pm$1.11
Avg. ACC 78.3$\pm$1.3 79.1$\pm$1.4 78.8$\pm$1.12 76.5$\pm$0.8 78.1$\pm$1.2 78.3$\pm$1.1
ECE 9.93$\pm$1.46 9.74$\pm$1.13 7.42$\pm$0.86 10.41$\pm$1.43 9.46$\pm$1.12 10.23$\pm$0.95
Table 2: Comparison of OOD accuracy and ECE across RMNIST, PACS, and VLCS datasets. Following hyperparameter tuning for each method, measurements were taken across three different random seeds, presenting the mean and variance for each metric. It was observed that on average, IB-IRM exhibits a lower ECE while maintaining comparable inference performance.
Figure 5: The relationship between out-of-distribution accuracy and ECE varies with the $\lambda$ parameter of IRMv1 in the RMNIST dataset. As the color shifts from blue to green, $\lambda$ increases, indicating stronger regularization in the model. Points closer to blue, situated on the middle right of the figure, correspond to intermediate ECE and high accuracy. In contrast, points nearer to green, located towards the lower left, indicate better ECE but lower accuracy. This suggests that in more practical datasets, such as those with diversity shifts, stronger regularization may improve ECE, but this often comes at the cost of a significant decrease in accuracy.
Considering that a model’s OOD accuracy affects its calibration (Ovadia et al., 2019), a comparative evaluation of ECE combined with OOD accuracy was conducted for a fairer comparison of ECE. A threshold was set for validation accuracy in the training environment, and upon reaching this threshold, early stopping of training was implemented to create conditions of comparable OOD accuracy.
The average results from three different random seeds under the aforementioned conditions are presented in Table 2, where although all methods achieved a similar average OOD accuracy, notable differences were observed in OOD ECE. Specifically, the ECE of IB-IRM was on average lower than that of any other method. Additionally, Figure 4 visualizes the relationship between ECE in the training and test environments, with the red solid line representing cases where ECE is equal in both environments. Ideally, models should be distributed near this red line from the perspective of cross-environmental calibration of uncertainty. Specifically, in Figures 3(a) and 3(c), IB-IRM is distributed near the red solid line compared to other methods, suggesting that it does not overfit to the training environment and exhibits more ideal feature learning.
It was also suggested that regularization in models such as IRMv1 and IBIRM, while reducing OOD ECE in cases of Diversity Shift, can adversely affect accuracy. Figure 5 illustrates the relationship between OOD accuracy and ECE in RMNIST when varying the magnitude of $\lambda$ in IRMv1. As regularization strengthens, the model’s ECE improves; however, this is accompanied by a drastic deterioration in accuracy, indicating a trade-off between the two.
6 Discussion and Conclusion
In this paper, focusing on IRM, which is challenging to realize due to the difficulties of bi-level optimization, we conducted comparative evaluations using ECE, a computationally feasible metric for calibration, to better understand its approximation methods.
The experimental results suggested that IB-IRM, which applies the information bottleneck method to IRM, is superior in terms of calibration. It is hypothesized that IB-IRM prevents excessively high confidence levels caused by dependence on environment-specific features by controlling the complexity of the feature extractor $\Phi(X)$. Additionally, it was suggested that regularization penalties introduced for the purpose of OOD generalization are effective in achieving lower OOD ECE. However, on the other hand, it was observed that in more practical datasets (Diversity Shift), while the ECE of the model becomes lower, the OOD accuracy worsens. This indicates a trade-off between OOD ECE and accuracy, suggesting that careful balancing of these two metrics is important in aiming for OOD generalization.
As a future challenge, it has become clear that the degree of invariant feature learning cannot be fully measured by accuracy in out-of-environment data alone, necessitating the establishment of broader generalization metrics and measurement methods. Furthermore, a better understanding of the relationship between OOD generalization and calibration, and the development of new approximation methods for IRM based on this understanding, are anticipated.
Acknowledgments
This work used computational resources TSUBAME3.0 supercomputer provided by Tokyo Institute of Technology through the TSUBAME Encouragement Program for Younger Users awarded to Kotaro Yoshida.